# Primordial black hole formation in $k$-inflation models

### v3

@@ 5.1 PBH mass fraction
where the height $\beta_{0}$ can be estimated using either (82) or (89). By way of example, we will use $\gamma=0.36$, $\delta_{c}=0.56$, and $\sigma=(4/9)\sqrt{\mathcal{P}_{0}/2}$ with $\mathcal{P}_{0}=0.2$ or $0.4$. Then, using (89) we obtain $\beta_{0}=1.536\cdot 10^{-7}$ for $\mathcal{P}_{0}=0.2$ and $3.429\cdot 10^{-5}$ for $\mathcal{P}_{0}=0.4$. In contrast, using (82) we obtain $\beta_{0}=3.441\cdot 10^{-5}$ and $2.489\cdot 10^{-3}$ for $\mathcal{P}_{0}=0.2$ and $\mathcal{P}_{0}=0.4$, respectively. Hence, the Press-Schechter prediction based on Eq. (82) is by a factor of 100 larger than the prediction of the CCP formalism based on Eq. (89). In the following, we will use the more conservative CCP formalism.
Nota Bene: The height $\beta_{0}$ decreases very quickly with decreasing $\mathcal{P}_{0}$. For example, using (89) we find for $\mathcal{P}_{0}=0.08$. For the values of $\mathcal{P}_{0}$ below 0.08, e.g., 0.04, the value of $\beta_{0}$ becomes insignificant.
@@ 5.2 PBH abundance
For an estimate based on the top-hat toy model, we use $\beta(M)$ defined by (90) with mass range $(M_{\rm min},M_{\rm max})$ defined by (78) and (79) and $\beta_{0}$ calculated in Sec. 5.1 for a chosen $\mathcal{P}_{0}$. It is clear that the values $\mathcal{P}_{0}$ and $q_{\rm max}$ (or the corresponding $M_{\rm min}$) are decisive since $\beta(M)$ is extremely sensitive to $\mathcal{P}_{0}$ and since $f_{\rm PBH}$ peaks just above $M_{\rm min}$.
@@ 6 Summary and conclusions
The PBH abundance is fairly sensitive to the value of the critical collapse threshold $\delta_{\rm c}$. We have found that the shape of the spectra in both models does not significantly change with slight variations in input parameters. As a consequence, the critical collapse threshold $\delta_{\rm c}$ is also not very sensitive to these variations. Hence, the values of $\delta_{\rm c}$ calculated for the spectra presented in Figs. 3 and 4 for the PLLS and the Tachyon model, respectively, could have been applied for each set of parameters presented in Table 2.
Our results for the PLLS model show that a suitable choice of input parameters different from those in Ref. [20] could yield a significant PBH abundance. For a chosen set of parameters yielding the power spectrum of Fig. 3, the PBH fraction has a few sharp peaks around the PBH mass $M=10^{-11}M_{\odot}$. In contrast, the Tachyon model power spectrum of Fig.  4 predicts experimentally allowed PBH abundance today in a broad mass range (Fig. 6). In the PLLS model, the integrated PBH abundance of the highest peak (full blue line) in Fig. 6 could account for total DM. In the Tachyon model, the highest integrated contribution of PBHs to DM estimate is about $11\%$ (full black line), except for a particular parameterization (long dashed black line) where the integrated PBH contribution may reach up to $98\%$. However, the latter result should not be taken seriously since the spectrum corresponding to this parameterization is not supported by the observational data near the CMB pivot scale.
@@ Acknowledgments
This work is supported by the ICTP-SEENET-MTP project NT-03 Cosmology-Classical and Quantum Challenges and the COST action CA1810 “Quantum gravity phenomenology in the multi-messenger approach”. N. Bilić is indebted to E. Saridakis for valuable comments. M. Stojanović acknowledges the support by the Ministry of Science, Technological Development and Innovation of the Republic of Serbia under contract 451-03-137/2025-03/200113. D.D. Dimitrijević, G.S. Djordjević, and M. Milošević acknowledge the support by the same Ministry under contract 451-03-137/2025-03/200124. G.S. Djordjević and and D.D. Dimitrijević acknowledge the support by the CEEPUS Program RS-1514-03-2223 “Gravitation and Cosmology”. G.S. Djordjević acknowledges the hospitality of the CERN-TH. N. Bilić acknowledges the hospitality of the Department of Physics, University of Niš, where a part of his work has been completed.

### v4

@@ 5.1.3 Comparison
where the height $\beta_{0}$ can be estimated using either (95) or (102). By way of example, we will use $\gamma=0.36$, $\delta_{c}=0.56$, and $\sigma=(4/9)\sqrt{\mathcal{P}_{0}/2}$ with $\mathcal{P}_{0}=0.2$ or $0.4$. Then, using (102) we obtain $\beta_{0}=1.536\cdot 10^{-7}$ for $\mathcal{P}_{0}=0.2$ and $3.429\cdot 10^{-5}$ for $\mathcal{P}_{0}=0.4$. In contrast, using (95) we obtain $\beta_{0}=3.441\cdot 10^{-5}$ and $2.489\cdot 10^{-3}$ for $\mathcal{P}_{0}=0.2$ and $\mathcal{P}_{0}=0.4$, respectively. Hence, for a broad power spectrum similar to the toy-model spectrum (78), one may expect that the PS prediction based on Eq. (95) will be by about a factor of 100 larger than the prediction of the CCP formalism based on Eq. (102).
For the reasons stated above, in the following, we will use the more conservative CCP formalism. However, it is fair to mention that the quantitative difference between the two methods could be reduced by an effective calibration — for example, by adjusting the collapse threshold $\delta_{\rm c}$ or redefining $\sigma$ so that PS and CCP results become more consistent. In our example above, the height of the mass fraction $\beta_{0}$ decreases very quickly with decreasing $\mathcal{P}_{0}$, to wit, with decreasing $\sigma$. Hence, using different window functions in the two approaches and thereby redefining $\sigma$ could reduce disagreement between PS and CCP methods. In a recent study [86], it has been shown that incorporating a smoothing procedure into the peak theory framework alleviates the tension between peaks theory and the Press–Schechter formalism.
@@ 5.2 PBH abundance
For an estimate based on the top-hat toy model, we use $\beta(M)$ defined by (103) with mass range $(M_{\rm min},M_{\rm max})$ defined by (91) and (92) and $\beta_{0}$ calculated in Sec. 5.1 for a chosen $\mathcal{P}_{0}$. Obviously, the values $\mathcal{P}_{0}$ and $q_{\rm max}$ (or the corresponding $M_{\rm min}$) are decisive since $\beta(M)$ is extremely sensitive to $\mathcal{P}_{0}$ and since $f_{\rm PBH}$ peaks just above $M_{\rm min}$.
@@
Figure 8: The mass fraction $\beta$ (left panel) and fraction of PBH dark matter (right panel) versus $M$ in the Tachyon model for various critical thresholds $\delta_{\rm c}$ and fixed set of parameters as in Table 1 with $\phi_{\rm in}=1.1M_{\rm Pl}$.
@@
5.2.1 Sensitivity to $\delta_{\rm c}$
The mass fraction $\beta$ and consequently the PBH abundance calculated in both PS and CCP formalisms is exponentially sensitive to the critical collapse threshold $\delta_{\rm c}$. The value of $\delta_{c}$ depends on the shape of the collapsing curvature power spectrum [90] (see also [10, 77, 92, 83, 91, 93, 94, 95, 96]). In appendix A, we outline the calculation procedure. For a broad power spectrum that could be approximated by (78), we recover $\delta_{c}\sim 0.56$ of Ref. [90]. For the PLLS and Tachyon models, the variations of input parameters and initial values in the interesting ranges do not substantially affect the shape of the spectra. The variations of the corresponding critical collapse thresholds are within $\delta_{\rm c}=0.51$ for the PLLS model and $\delta_{\rm c}=0.56$ for the Tachyon model spectra.
In Fig. 8 we plot the mass fraction $\beta$ and the PBH abundance $f_{\rm PBH}$ in the Tachyon model for various $\delta_{c}$ within the plausible uncertainty departure of about 1% from the numerical value calculated in appendix A.
5.3 Comment on primordial non-Gaussianity
The three-point correlation function
$\langle\hat{\zeta}_{q_{1}}\hat{\zeta}_{q_{2}}\hat{\zeta}_{q_{3}}\rangle=(2\pi)^{3}\delta(\mbox{$q$}_{1}+\mbox{$q$}_{2}+\mbox{$q$}_{3})f_{\rm NL}F(q_{1},q_{2},q_{3})$ (110)
provides the prime diagnostic of non-Gaussianity of inflationary fluctuations [97]. The quantity $f_{\rm NL}$ is a dimensionless parameter defining the amplitude of non-Gaussianity and the function $F$ captures the momentum dependence.
The lowest-order non-Gaussian corrections to the primordial power spectrum arise from a single one-loop diagram, which leads to [98, 99, 100]
$\mathcal{P}_{\rm S}(q)=\mathcal{P}_{\rm S,G}(q)+f_{\rm NL}^{2}q^{3}\int\frac{d^{3}k}{4\pi}\frac{\mathcal{P}_{\rm S,G}(k)}{k^{3}}\frac{\mathcal{P}_{\rm S,G}(|\mbox{$q$}-\mbox{$k$}|)}{|\mbox{$q$}-\mbox{$k$}|^{3}},$ (111)
where $\mathcal{P}_{\rm S,G}$ is the dimensionless power spectrum defined in (50), and the second term on the right-hand side is a non-Gaussian contribution which can be rewritten as
$\mathcal{P}_{\rm S,NG}(q)=f_{\rm NL}^{2}q^{2}\int_{0}^{\infty}\frac{dk}{k^{2}}\mathcal{P}_{\rm S,G}(k)\int_{|q-k|}^{q+k}\frac{dk^{\prime}}{{k^{\prime}}^{2}}\mathcal{P}_{\rm S,G}(k^{\prime}).$ (112)
Ünal [99] estimated (112) for the log-normal gaussian spectrum
$\mathcal{P}_{\rm S,G}=A\exp[-\ln^{2}(k/k_{\rm peak})/(2\sigma^{2})].$ (113)
For an amplitude $A=0.01$, and $f_{\rm NL}=3$, he found that the non-Gaussian term is subdominant by about a factor of 10. Clearly, if the amplitude were 10 times larger, the non-Gaussian term would be of the same order as the Gaussian one. Similarly, Riccardi, Taoso, and Urbano [101] found that in realistic single-field inflationary models with ultra slow-roll, by changing the peak amplitude of the curvature power spectrum by about a factor of two, one could obtain the same abundance calculated with the Gaussian approximation as the one obtained in the presence of local non-Gaussianities.
In the $k$-essence models, the largest non-Gaussianity peakes at the so-called equilateral configuration with $q_{1}\sim q_{2}\sim q_{3}$. The non-Gaussianity amplitude of an equilateral triangle in a general $k$-essence is given by [97, 102]
$f_{\rm NL}^{\rm equil}=-\frac{35}{108}\left(\frac{1-c_{\rm s}^{2}}{c_{\rm s}^{2}}\right)+\frac{20}{81}\Lambda,$ (114)
where
$\Lambda\equiv X^{2}\frac{\mathcal{L}_{XX}^{2}-(1/3)\mathcal{L}_{X}\mathcal{L}_{XXX}}{\mathcal{L}_{X}^{2}+2X\mathcal{L}_{X}\mathcal{L}_{XX}}.$ (115)
The quantity $\Lambda$ in the Tachyon model is identically zero. As the sound speed deviates from unity most at the end of the slow roll regime, we estimate the equilateral amplitude at the end of inflation, neglecting possible post-inflationary effects. For an estimate we can use $c_{\rm s}^{2}=1/3$, given by (43) for $\varepsilon_{1}=1$, yielding
$f_{\rm NL}^{\rm Tachyon}\approx\left.f_{\rm NL}^{\rm equil}\right|_{\varepsilon_{1}=1}=-\frac{70}{108}.$ (116)
In the PLLS model, we find
$\Lambda=\frac{1-3\alpha+2\alpha^{2}}{3+6\alpha},$ (117)
so for $\alpha=1.5$ we have $\Lambda=1/12$. Thus, using (114) with $c_{\rm s}^{2}=1/2$, for the PLLS model we find $f_{\rm NL}\approx f_{\rm NL}^{\rm equil}=-295/972$
$f_{\rm NL}^{\rm PLLS}\approx\left.f_{\rm NL}^{\rm equil}\right|_{\varepsilon_{1}=1}=-\frac{295}{972}.$ (118)
Both values in (116) and (118) are well within the observational constraints provided by the Planck 2018 collaboration [103]: $f_{\rm NL}^{\rm equil}=-26\pm 47$.
To estimate the NG contribution to the power spectrum in the PLLS model, we could approximate the spectrum of Fig. 4 by the log-normal function (113) with an amplitude of about $A=0.1$. In our estimate, $f_{\rm NL}\simeq 0.3$, so the NG contribution to the power spectrum would amount to a factor of $(0.3/3)^{2}=0.01$ suppression compared to the NG term in Ref. [99]. Since our amplitude is a factor of 10 larger than that of Ref. [99], the NG contribution enhances the PLLS power spectrum of Fig. 4 by about 10%.
The results obtained for the log-normal spectrum of Ref. [99] cannot be used for the broad spectrum of our Tachyon model. However, we can obtain a rough estimate by making use of the top-hat power spectrum (78)). In this model, we find
$\mathcal{P}_{\rm S,NG}=f_{\rm NL}^{2}\mathcal{P}_{0}^{2}q^{2}\int_{q_{\rm min}}^{q_{\rm max}}\frac{dk}{k^{2}}\left[\frac{\Theta(|q-k|-q_{\rm min})\Theta(q_{\rm max}-|q-k|)}{|q-k|}\right$.
$\left.-\frac{\Theta(q+k-q_{\rm min})\Theta(q_{\rm max}-q-k)}{q+k}\right].$ (119)
The integration yields
$\mathcal{P}_{\rm S,NG}=f_{\rm NL}^{2}\mathcal{P}_{0}^{2}\left[2\ln\frac{q^{2}-q_{\rm min}^{2}}{q_{\rm min}^{2}}+2\ln\frac{q_{\rm max}-q}{q_{\rm max}}-\frac{2q^{2}}{q^{2}-q_{\rm min}^{2}}+\frac{q}{q_{\rm max}}+\frac{q}{q_{\rm max}-q}\right].$ (120)
In the middle of the distribution, where $q_{\rm min}\ll q\ll q_{\rm max}$ we find an estimate
$\mathcal{P}_{\rm S,NG}\simeq 2f_{\rm NL}^{2}\mathcal{P}_{0}^{2}\left(\ln\frac{q^{2}}{q_{\rm min}^{2}}-1+\mathcal{O}(q_{\rm min}^{2}/q^{2})+\mathcal{O}(q^{2}/q_{\rm max}^{2})\right).$ (121)
For our estimated $f_{\rm NL}\simeq-\frac{70}{108}$ and the parameters of our top-hat model of Fig. 6 $\mathcal{P}_{0}=0.09$, $q/q_{\rm min}=10^{4}\div 10^{8}$, we find $\mathcal{P}_{\rm S,NG}=0.12\div 0.24$, to be compared with $\mathcal{P}_{\rm S,G}=0.09$. Hence, the NG contribution in the top-hat model is positive and of the order $\mathcal{O}(1\div 2)\mathcal{P}_{\rm S,G}$. Thus, we expect the peak of the power spectrum that includes a non-Gaussian contribution in the Tachyon model to be enhanced by about a factor of $\mathcal{O}(2\div 3)$ compared with the Gaussian power spectrum.
However, there is in addition an intrinsic non-Gaussianity owing to the nonlinear relation (70) between $\zeta$ and $\delta_{r}$ [81, 104, 105]. In the presence of such non-Gaussianity, the amplitude of the power spectrum can be significantly reduced by a factor $\mathcal{O}(2\div 3)$ [105] compared to the Gaussian (linear) case, where one approximates (70) by the linear relation (71). In this way, the effect of this intrinsic non-Gaussianity works in the opposite direction compared to the primordial non-Gaussianity discussed above, so these two effects might partly annihilate. However, a precise estimate of the total outcome requires more accurate calculations, which go beyond the scope of the present paper.
Approximating the nonlinear relation (70) by the linear one can also affect the critical threshold $\delta_{\rm c}$. According to Kehagias, Musco, and Riotto [106], the impact of the intrinsic non-Gaussianity on a relative change of $\delta_{\rm c}$ with respect to the value obtained using a linear relation between $\zeta$ and $\delta_{r}$ is of the order of a few percent.
@@ 6 Summary and conclusions
The PBH abundance is fairly sensitive to the value of the critical collapse threshold $\delta_{\rm c}$. We have found that the shape of the spectra in both models does not significantly change with slight variations in input parameters. As a consequence, the critical collapse threshold $\delta_{\rm c}$ is also not very sensitive to these variations. Hence, the values of $\delta_{\rm c}$ calculated for the spectra presented in Figs. 4 and 5 for the PLLS and the Tachyon model, respectively, could have been applied for each set of parameters presented in Table 2. However, as Fig. 8 shows, the relative change of $\delta_{\rm c}$ (e.g., caused by the intrinsic non-Gaussianity [106]) within about $\pm$1% could affect the PBH production by a factor of up to $4\div 6$.
Our results for the PLLS model show that a suitable choice of input parameters different from those in Ref. [31] could yield a significant PBH abundance. For a chosen set of parameters yielding the power spectrum of Fig. 4, the PBH fraction has a few sharp peaks around the PBH mass $M=10^{-11}M_{\odot}$. In contrast, the Tachyon model power spectrum of Fig. 5 predicts experimentally allowed PBH abundance today in a broad mass range (Fig. 7). In the PLLS model, the integrated PBH abundance of the highest peak (full blue line) in Fig. 7 could account for total DM. In the Tachyon model, the highest integrated contribution of PBHs to DM estimate is about $11\%$ (full black line), except for a particular parameterization (long dashed black line) where the integrated PBH contribution may reach up to $98\%$. However, the latter result should not be taken seriously since the spectrum corresponding to this parameterization is not supported by the observational data near the CMB pivot scale.
@@ Acknowledgments
This work is supported by the ICTP-SEENET-MTP project NT-03 Cosmology-Classical and Quantum Challenges and the COST action CA1810 “Quantum gravity phenomenology in the multi-messenger approach”. The work of N. Bilić is supported by the COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), while G. S. Djordjević acknowledges support by the COST Action CA23130 - Bridging high and low energies in search of quantum gravity (BridgeQG). N. Bilić is indebted to E. Saridakis for valuable comments. M. Stojanović acknowledges the support by the Ministry of Science, Technological Development and Innovation of the Republic of Serbia under contract 451-03-137/2025-03/200113. D.D. Dimitrijević, G.S. Djordjević, and M. Milošević acknowledge the support by the same Ministry under contract 451-03-137/2025-03/200124. G.S. Djordjević and and D.D. Dimitrijević acknowledge the support by the CEEPUS Program RS-1514-03-2223 “Gravitation and Cosmology”. G.S. Djordjević acknowledges the hospitality of the CERN-TH. N. Bilić acknowledges the hospitality of the Department of Physics, University of Niš, where a part of his work has been completed.